# The Tautochrone of Huygens and Abel: From Constructive Geometry to Fractional Calculus

### v1

@@ 3.3 The Birth of Fractional Calculus
Let us pause for a moment to underline a central point of the present analysis. We remember that the integral equation obtained by Abel in the tautochrone problem can formally be written as
$\int_{0}^{x}\frac{g(t)}{\sqrt{(x-t)^{1/2}}}dt=f(x)$ (27)
However, this is, essentially, the definition of a fractional integral of order $1/2$, in the sense of Riemann–Liouvile, as we discuss soon. We may rewrite it in the form:
$_{0}{\textrm{I}}_{x}^{1/2}g(x)=\int_{0}^{x}\frac{g(t)}{\sqrt{(x-t)^{1/2}}}dt,$ (28)
where we have introduced the notation $_{0}{\textrm{I}}_{x}^{1/2}$ for an integration of arbitrary order $1/2$ in the variable $x$. Thus, the tautochrone problem in the Abel’s perspective may assume the form:
$f(x)={}_{0}{\textrm{I}}_{x}^{1/2}g(x).$ (29)
In this way as well, to obtain the unknown function $g(x)$ became possible if we promote the inversion of a fractional operator, i.e., of the fractional integral of order $1/2$. In this framework, the Abel problem would be the particular case of the inversion of a fractional operator. The formal solution would require the application of the corresponding fractional derivative, i.e.,
$g(x)=\frac{d^{1/2}}{dx^{1/2}}\int_{0}^{x}\frac{f(t)}{(x-t)^{1/2}}dt.$ (30)
This kind of simplified reasoning opens space to discuss the fractional calculus as a natural generalization of the classical differential operators, motivated by historical problems. Thus, the modern interpretation of the Abel problem as involving a fractional operator is an instructive bridge between the early years of the mechanics and mathematical analysis and the present.
Let us now move to a more formal way to connect again these results from Abel to the roots of the fractional calculus. To accomplish this task, the problem can be reformulated in such a way that one has to find $f(z)=ds/dz$, given $k(y)=k$, such that:
$k(y)=k=\int_{0}^{y}\frac{f(z)}{(y-z)^{1/2}}dz.$ (31)
Now, we proceed in a different manner, multiplying both sides of (31) by $1/\Gamma(1/2)=1/\sqrt{\pi}$ in order to obtain
$\frac{k}{\Gamma(1/2)}$ $=$ $\frac{1}{\Gamma(1/2)}\int_{0}^{y}\frac{f(z)}{(y-z)^{1/2}}dz$ (32)
$=$ $_{0}{\textrm{I}}_{z}^{1/2}\left[f(z)\right],$
whose right-hand side we can recognize (better: we define!) as a semi-integral operator, that is, a fractional integral operator of order $1/2$, represented by the symbol $_{0}{\textrm{I}}_{z}^{1/2}$ we have introduced before. If this is true, to cancel the “semi-operator”, we have just to take the “semi-derivative” of each side of (32), that is,
$\frac{d^{1/2}}{dz^{1/2}}\left[\frac{k}{\Gamma(\alpha)}\right]=\frac{d^{1/2}}{dz^{1/2}}\bigg{\{}{}_{0}{\textrm{I}}_{z}^{1/2}\left[f(z)\right]\bigg{\}}=f(z).$ (33)
In this way, the problem is reduced to calculate the derivative of order $1/2$ of the left-hand side, i.e., of a constant. This calculation may be performed by using a result due to Lacroix [27].
Indeed, a derivative of arbitrary order was briefly mentioned (perhaps for the first time as such) in 1819, in the book of Sylvestre François Lacroix (1765–1843), from which an explicit formula for a fractional derivative may be obtained. The central argument may be summarized as follows (using a today’s notation). Let, for instance, $y=x^{m}$; when $n$ is an integer, one has for any arbitrary $m\in\mathbb{N}$:
$d^{n}y=d^{n}(x^{m})$ $=$ $m(m-1)\cdots(m-n+1)x^{m-n}dx^{n}$ (34)
$=$ $\frac{\Gamma(m+1)}{\Gamma(m-n+1)}x^{m-n}dx^{n},\quad n\in\mathbb{N},$
where $\mathbb{N}=\{1,2,3,$...\} is the set of natural numbers. If we put $m=2$ and $n=1$, then we obtain
$d^{1}(x^{2})=2x\,dx,$ (35)
as expected. Now, we may consider $m=1$ and $n=1/2$ and try to answer to the original question. We obtain:
$d^{1/2}x=\frac{\Gamma(2)}{\Gamma(3/2)}x^{1/2}dx^{1/2}=\frac{2}{\sqrt{\pi}}\sqrt{x\,dx}.$ (36)
Thus, the result obtained by Lacroix may be put in the usual form as:
$\frac{d^{1/2}x}{dx^{1/2}}=\frac{2\sqrt{x}}{\sqrt{\pi}},$ (37)
which, apart from a constant, may be compared with the result reported by Leibniz, in his letter dated September 30, 1695, to L’Hôpital [28]:
You see from this, Sir, that one can express by an infinite series a quantity such as $d^{1/2}\overline{xy}$, or $d^{1:2}\overline{xy}$, even though this may seem far removed from Geometry, which ordinarily knows only differences with positive integer exponents, or negative ones with respect to sums, and not yet those whose exponents are fractional. It is true that it remains to provide $d^{1:2}x$ for that series; but even this can be explained in some fashion. For let the ordinates $x$ be in geometric progression, so that, taking a constant $d\beta$, one has $dx=xd\beta:a$, or (taking $a$ as unity) $dx=xd\beta$. Then ddxddx will be $x\overline{d\beta}^{2}$, and $dx^{3}$ will be $x\overline{d\beta}^{3}$, etc., and $d^{e}x=x\cdot\overline{d\beta}^{e}$. And by this device, the potential exponent, and replacing $dx:x$ with $d\beta$, we have $d^{e}x=\overline{dx:x}^{e}\cdot x$. Thus it follows that $d^{1:2}x$ will be equal to $x\cdot\sqrt[2]{dx:x}$. It seems likely that one day useful consequences will be drawn from these seemingly useless paradoxes. You are among those who can go furthest in discoveries, and I shall soon be obliged to hand over the torch to others. I wish I had much to share, for this verse: your knowledge is nothing unless someone else knows that you know is most true, in that thoughts which were of little value in themselves can give rise to far more beautiful ones.
Thus, the result presented by Leibniz for the question raised by L’Hôpital may be interpreted as stating that for the fractional exponent $e$, one has:
$d^{e}x=x^{1-e}dx^{e},$ (38)
which could be compared with Eq. (34), because, for $m=1$ and $n=e$, it could be rewritten as:
$d^{e}x=\frac{\Gamma(2)}{\Gamma(2-e)}x^{1-e}dx^{e}.$ (39)
In this framework, if we consider consider the case $n=e=1/2$ and $m=0$, that is, $y=x^{0}=1$, from Eq. (34), we obtain the “surprising” result that the fractional derivative of a constant is not zero, but instead a function of $x$, namely:
$\frac{d^{1/2}}{dx^{1/2}}1=\frac{1}{\sqrt{\pi x}}.$ (40)
Coming back to our problem, by using the results above in Eq. (33), we obtain:
$f(z)=\frac{k}{\pi\sqrt{z}}.$ (41)
This permits us to deduce that the answer to the tautochrone problem is formally:
$\frac{ds}{dy}=\frac{\sqrt{2g}}{\pi}\frac{T(y_{0})}{\sqrt{y}}.$ (42)
Integrating the above equation, using the condition $s(0)=0$, we obtain:
$s(y)=2\sqrt{2g\left[\frac{T(y_{0})}{\pi}\right]^{2}}y^{1/2},$ (43)
which is the cycloid sketched in Fig. 3, as shown by Abel in connection with this problem of finding the path followed by the particle in order the time of sliding down be independent of the initial height [17].
In summary, to solve the tautochrone problem, in the pioneering Abel’s perspective, we have to determine $f(z)$, a task that can be accomplished by computing the fractional derivative of a constant $k$. In the paper of Abel, it is not clear that if he used the Lacroix or its own results. Anyway, in the paper of 1826, Abel provided the solution for the integral equation [18]:
$k(y)=\int_{c}^{y}\frac{f(z)}{(y-z)^{\alpha}}dz,\quad y>c,\quad 0<\alpha<1.$ (44)
In practice, Abel used the operators that nowadays are ascribed to Riemann and Liouville, preceding them by at least one decade [29, 30]. To complete the analysis, let us formalize a little bit more the previous results, considering the Abel fractional integral equation of first kind, defined as:
$_{0}{\textrm{I}}_{t}^{\alpha}g(t)=\frac{1}{\Gamma(\alpha)}\int_{0}^{t}\frac{g(\tau)}{(t-\tau)^{1-\alpha}}d\tau=f(t),\quad 0<\alpha<1.$ (45)
Equation (21) is a particular case of Eq. (45), for $\alpha=1/2$. As we have seen above, Eq. (45) may be solved in terms of a fractional derivative in the form:
$_{0}{\textrm{D}}_{t}^{\alpha}\left[{}_{0}{\textrm{I}}_{t}^{\alpha}g(t)\right]=g(t)={}_{0}{\textrm{D}}_{t}^{\alpha}f(t),$ (46)
where we have also introduced the symbol $_{0}{\textrm{D}}_{t}^{\alpha}$ now to denote a derivative of arbitrary order $\alpha$ in the variable $t$. A notation of this kind was introduced by the mathematician Harold Thayer Davis (1892–1974) [31] and is useful to highlight the meaning of the left subscript, which is not obvious in a derivative [32].
In Eq. (46), we have supposed that $_{0}{\textrm{D}}_{t}^{\alpha}$ exists (even if we have not discovered its form until now!) and is the left-inverse of the operator $_{0}{\textrm{I}}_{t}^{\alpha}$, that is,
$_{0}{\textrm{D}}_{t}^{\alpha}{}_{0}{\textrm{I}}_{t}^{\alpha}=\mathds{1},$ (47)
where $\mathds{1}$ is the identity operator. A solution for (45) may be now searched by using the Laplace transform and the convolution theorem [32, 33].
4 The Fractional Operators of Riemmann-Liouville and Caputo
The expression fractional calculus we have been using until now — although something of a misnomer — refers to the theory concerning integrals and derivatives of arbitrary order. In this context, it may be viewed as a natural extension of the classical concepts of differentiation and integration to non-integer (fractional) orders. A conventional way to introduce the basic ideas of fractional calculus is by posing a question that has become emblematic: what meaning can we assign to the expression
$\frac{d^{1/2}f(x)}{dx^{1/2}}?$ (48)
This very question was famously raised by Guillaume François Antoine, marquis de L’Hôpital (1661–1704), in a query addressed to Gottfried Wilhelm Leibniz (1646–1716), the founder of differential calculus, whose answer we quoted after Eq. (36) in discussing the Lacroix derivative [28].
The foundational questions that marked the birth of fractional calculus — once seen as paradoxical or speculative — have now been thoroughly examined and answered in full detail. These inquiries reveal profound conceptual challenges. For instance, if the first derivative of a function represents its instantaneous rate of change or the slope at a point, what geometric meaning can we assign to a derivative of order $1/2$? And if we persist in exploring the notion of half-order differentiation, we are naturally led to ask: what operator, when applied twice, yields the first derivative? Such questions, once open-ended, have been rigorously addressed, and the field of fractional calculus has since matured into a well-established and robust branch of mathematical analysis [34, 35, 36].
Today, the Riemann-Liouville fractional integral or Riemann-Liouville integral of arbitrary order $\alpha$ is defined as:
$_{c}{\textrm{D}}_{x}^{-\alpha}f(x)=\frac{1}{\Gamma(\alpha)}\int_{c}^{x}\frac{f(t)}{(x-t)^{1-\alpha}}dt,\quad\quad\Re(\alpha)>0,$ (49)
where $\Re(\dots)$ denotes the real part of $\dots$. Then, it is clear that the fractional derivative of arbitrary order $\alpha$ may be interpreted as the derivative of positive integer order of a derivative of negative non-integer order, which is an integral of arbitrary non-integer order. Accordingly, the operator
${}_{c}{\textrm{D}}_{x}^{\alpha}f(x)=\frac{1}{\Gamma(k-\alpha)}\frac{d^{k}}{dx^{k}}\left[\int_{c}^{x}\frac{f(t)}{(x-t)^{\alpha+1-k}}dt\right],\quad k\in\mathbb{N},$ (50)
for $\alpha=k-p$ is the definition of the Riemann-Liouville fractional derivative to be used hereafter.
A concise, alternative way, to face the problem of using the definition of an integral of arbitrary order to obtain a definition of a derivative of arbitrary order may be sketched as follows. We start with the Riemann-Liouville integral defined from the Cauchy integral formula, as before:
$_{c}{\textrm{I}}_{x}^{\alpha}f(x)=\frac{1}{\Gamma(\alpha)}\int_{c}^{x}(x-t)^{\alpha-1}f(t)dt.$ (51)
Notice that $\alpha$ may be complex, with the real part strictly positive, that is, $\Re(\alpha)>0$. We assume for a moment that $\alpha\in\mathbb{R}$. The definition (51) has the following properties:
${}_{c}{\textrm{I}}_{x}^{\alpha}\left[{}_{c}{\textrm{I}}_{x}^{\beta}f(x)\right]={}_{c}{\textrm{I}}_{x}^{\alpha+\beta}f(x)\quad{\textrm{and}}\quad\frac{d}{dx}\left[{}_{c}{\textrm{I}}_{x}^{\alpha+1}f(x)\right]={}_{c}{\textrm{I}}_{x}^{\alpha}f(x).$
One is then tempted to assume that the fractional derivative could be easily defined as
$_{c}{\textrm{D}}_{x}^{\alpha}f(x)={}_{c}{\textrm{I}}_{x}^{-\alpha}f(x).$ (52)
One of the problems with this identification is that the gamma function is not defined for zero or negative integers. Anyway, we may follow a procedure inspired in this idea with some caution. We notice that for $n\in\mathbb{N}$, we surely write:
$\frac{d^{n}}{dx^{n}}\left[{}_{c}{\textrm{I}}_{x}^{n}f(x)\right]=f(x),$ (53)
that is, taking $n$ times the derivative of a function $f(x)$ after integration it $n$ times is equivalent to the identity operator, as stated in (47). With this result in mind, we expect to find a fractional derivative operator such that
$_{c}{\textrm{D}}_{x}^{\alpha}\left[{}_{c}{\textrm{I}}_{x}^{\alpha}f(x)\right]=f(x).$ (54)
An operator that can be constructed is
$_{c}{\textrm{D}}_{x}^{\alpha}f(x)=\frac{d^{\lceil{\alpha}\rceil}}{dx^{\lceil{\alpha}\rceil}}\left[{}_{c}{\textrm{I}}_{x}^{{\lceil{\alpha}\rceil}-\alpha}f(x)\right],$ (55)
with $\lceil{\alpha}\rceil$} being the ceiling function, which gives the smallest integer greater or equal to $\alpha$ (that is, one has to consider the next integer). This permits us to arrive at the operator defined in (50) for $\alpha=k-p$, such that $0<p=k-\alpha<1$, and $_{c}{\textrm{D}}_{x}^{k}=d^{k}/dx^{k}$ being the usual derivative of integer order. It is, indeed, the left Riemann-Liouville fractional derivative [37].
Along the same lines, the Caputo operator may be defined as follows [24, 25]:
$^{\textrm{C}}_{c}{\textrm{D}}_{x}^{\alpha}f(x)=\frac{1}{\Gamma(m-\alpha)}\int_{c}^{x}\frac{f^{(m)}(t)}{(x-t)^{\alpha+1-m}}dt,\quad\quad m-1<\alpha<m,$ (56)
which, for $m=1$, becomes:
$^{\textrm{C}}_{c}{\textrm{D}}_{x}^{\alpha}f(x)=\frac{1}{\Gamma(1-\alpha)}\int_{c}^{x}\frac{f^{\prime}(t)}{(x-t)^{\alpha}}dt,\quad\quad 0<\alpha<1.$ (57)
Apart the constant coefficient represented by $1/\Gamma(1-\alpha)$, the operator defined in Eq. (57) coincides with the definition used by Abel — the one we have rewritten in Eq. (16). To conclude the abbreviated presentation of these operators, we remark that a relation between the Caputo and the Riemann-Liouville derivatives may be established as follows [35]:
$^{\textrm{C}}_{0}{\textrm{D}}_{x}^{\alpha}f(x)={}_{0}{\textrm{D}}_{x}^{\alpha}f(x)-\sum_{k=0}^{n-1}\frac{f^{(k)}(0)x^{k-\alpha}}{\Gamma(k-\alpha+1)},\quad n-1<\alpha<n,$ (58)
which, when $0<\alpha<1$, yields
${}^{\textrm{C}}_{0}{\textrm{D}}_{x}^{\alpha}f(x)$ $=$ ${}_{0}{\textrm{D}}_{x}^{\alpha}\left[f(x)-f(0)\right]$ (59)
$=$ ${}_{0}{\textrm{D}}_{x}^{\alpha}f(x)-\frac{f(0)x^{-\alpha}}{\Gamma(1-\alpha)}.$
We notice that, from the above equation, Caputo and Riemann-Liouville operators coincide for the functions such that $f(0)=0$. In addition, if we put now $c=$-$\infty$ in the Riemann-Liouville derivative, then it becomes
${}_{-\infty}{\textrm{D}}_{x}^{\alpha}f(x)$ $=$ $\frac{d^{m}}{dx^{m}}\left[\frac{1}{\Gamma(m-\alpha)}\int_{-\infty}^{x}\frac{f(t)}{(x-t)^{\alpha+1-m}}dt\right]$ (60)
$=$ $\frac{1}{\Gamma(m-\alpha)}\int_{-\infty}^{x}\frac{f^{(m)}(t)}{(x-t)^{\alpha+1-m}}dt$
$=$ ${{}_{-\infty}\!\!}{}^{\textrm{C}}{\textrm{D}}_{x}^{\alpha}f(x),\quad\quad m-1<\alpha<m.$
In this limit, both definitions become equal provided that $f(x)$ and its derivatives have a reasonable behavior when $x\to\infty$, that is, $f^{(k)}(-\infty)\to 0$, for k=0,1,,n1k=0,1,...,n-1, with $n={\lceil{\alpha}\rceil}$. This is a very important property from the physical point of view because to consider stationary processes is now permitted as, for instance, in the response of the fractional order dynamical systems to a periodic signal, required in the impedance problems [38], in the wave propagation in continuous media [32, 39], etc.
5 Concluding Remarks

### v2

@@ III.3 The Birth of Fractional Calculus
Let us pause for a moment to underline a central point of the present analysis. Abel’s solution to the tautochrone problem marks a pivotal moment in the history of mathematics: the emergence of fractional calculus. Equation (11),
$k=\int_{0}^{y}\frac{f(z)}{\sqrt{y-z}}\mskip 3.0mudz,$
is not merely an integral equation; it defines a fractional integral of order $1/2$. Introducing the notation Davis :
$_{0}\mathrm{I}_{y}^{1/2}f(y)=\frac{1}{\Gamma(1/2)}\int_{0}^{y}\frac{f(z)}{\sqrt{y-z}}\mskip 3.0mudz,$
the tautochrone condition becomes $_{0}\mathrm{I}_{y}^{1/2}f(y)=\text{constant}$. Solving it requires inverting this fractional operator — i.e., applying a fractional derivative $d^{1/2}/dy^{1/2}$.
Abel’s Insight: Inversion via Fractional Derivatives.
Abel recognized that if
$k={}_{0}\mathrm{I}_{y}^{1/2}f(y),$
then formally
$f(y)=\frac{d^{1/2}}{dy^{1/2}}\mskip 3.0muk.$
Thus, the problem reduces to computing the fractional derivative of a constant. In his 1826 paper Abel1826 , Abel solved the more general equation
$k(y)=\int_{c}^{y}\frac{f(z)}{(y-z)^{\alpha}}\mskip 3.0mudz,\qquad 0<\alpha<1,$ (14)
thereby introducing what would later be called the Riemann–Liouville fractional integral. His solution,
$f(y)=\frac{d^{\alpha}}{dy^{\alpha}}\mskip 3.0muk(y),$
treats fractional integration and differentiation as inverse operations — a foundational idea of fractional calculus.
Early Fractional Derivative: Lacroix and Leibniz.
The computation of $d^{1/2}(1)/dy^{1/2}$ had precursors. In 1819, Lacroix noted that for $y=x^{m}$,
$\frac{d^{n}}{dx^{n}}x^{m}=\frac{\Gamma(m+1)}{\Gamma(m-n+1)}x^{m-n},\quad n\in\mathbb{N},$
and extended the formula to non-integer $n$. Setting $m=0$, $n=1/2$ gives
$\frac{d^{1/2}}{dx^{1/2}}1=\frac{1}{\sqrt{\pi x}},$
showing that the fractional derivative of a constant is non-zero Lacroix1819 . This result echoes Leibniz’s 1695 speculation that $d^{1/2}x=x\sqrt{dx/x}$ Leibniz1695 — an early hint that differentiation could admit fractional orders.
Fractional Calculus as a Unified Framework.
Abel’s work implicitly defined the Abel integral equation of the first kind:
$_{0}\mathrm{I}_{t}^{\alpha}g(t)=\frac{1}{\Gamma(\alpha)}\int_{0}^{t}\frac{g(\tau)}{(t-\tau)^{1-\alpha}}\mskip 3.0mud\tau=f(t),\quad 0<\alpha<1.$ (15)
Its solution,
$g(t)={}_{0}\mathrm{D}_{t}^{\alpha}f(t),$
assumes the existence of a left–inverse operator satisfying
$_{0}\mathrm{D}_{t}^{\alpha}\mskip 3.0mu{}_{0}\mathrm{I}_{t}^{\alpha}=\mathds{1}.$
This operator — the Riemann–Liouville fractional derivative — was systematized later, but Abel had already used it to solve the tautochrone problem.
Thus, the tautochrone problem served as the first concrete physical motivation for fractional calculus. Abel’s analysis transformed a geometric question (Huygens) into an analytic one, inaugurating a new branch of mathematics that would later find applications across physics, engineering, and applied analysis Mainardi2007 ; Mainardi2007b ; Davis ; Mainardi-Book ; Evangelista2023 ; Oldham1974 ; Podlubny ; Samko ; Yang-Gao-Ju .
IV Fractional Operators: From Riemann–Liouville to Caputo
Fractional calculus extends differentiation and integration to arbitrary (non-integer) orders Lacroix1819 . The emblematic question — what meaning can be given to $d^{1/2}f/dx^{1/2}?$ — was first posed by L’Hôpital to Leibniz in 1695 Leibniz1695 . Today, fractional operators are rigorously defined and widely used in applications.
Riemann–Liouville Operators.
The Riemann–Liouville fractional integral of order $\alpha>0$ is
$_{c}\mathrm{I}_{x}^{\alpha}f(x)=\frac{1}{\Gamma(\alpha)}\int_{c}^{x}\frac{f(t)}{(x-t)^{1-\alpha}}\mskip 3.0mudt.$ (16)
For $\alpha\in(m-1,m)$, the corresponding fractional derivative is defined as
$_{c}\mathrm{D}_{x}^{\alpha}f(x)=\frac{d^{m}}{dx^{m}}\left[{}_{c}\mathrm{I}_{x}^{m-\alpha}f(x)\right],$ (17)
ensuring that $_{c}\mathrm{D}_{x}^{\alpha}\mskip 3.0mu{}_{c}\mathrm{I}_{x}^{\alpha}=\mathds{1}$. These operators satisfy the expected composition laws and reduce to ordinary calculus when $\alpha$ is an integer.
Caputo Operators.
In 1967, Caputo introduced an alternative definition that interchanges the order of differentiation and integration Caputo1967 ; Caputo1969 :
$_{c}^{\mathrm{C}}\mathrm{D}_{x}^{\alpha}f(x)={}_{c}\mathrm{I}_{x}^{m-\alpha}\left[\frac{d^{m}}{dx^{m}}f(x)\right],\qquad m-1<\alpha<m.$ (18)
For $0<\alpha<1$, this becomes
$_{c}^{\mathrm{C}}\mathrm{D}_{x}^{\alpha}f(x)=\frac{1}{\Gamma(1-\alpha)}\int_{c}^{x}\frac{f^{\prime}(t)}{(x-t)^{\alpha}}\mskip 3.0mudt.$ (19)
Apart from the normalizing factor $1/\Gamma(1-\alpha)$, Eq. (19) is precisely the operator that appears in Abel’s solution of the tautochrone problem (cf. Eq. (9)). Indeed, Abel’s integral equation
$k=\int_{0}^{y}\frac{f(z)}{(y-z)^{1/2}}\mskip 3.0mudz$
is, up to a constant, a Caputo derivative of order $1/2$.
Relation Between the Two Formulations.
The Riemann–Liouville and Caputo derivatives are connected by
$_{c}^{\mathrm{C}}\mathrm{D}_{x}^{\alpha}f(x)={}_{c}\mathrm{D}_{x}^{\alpha}\bigl[f(x)-f(c)\bigr],\qquad 0<\alpha<1.$ (20)
Thus they coincide when $f(c)=0$. For $c\to$-$\infty$, both definitions also coincide provided the function and its derivatives vanish at infinity — a property crucial for modelling stationary processes in fractional dynamical systems, wave propagation, and impedance problems Evangelista2018 ; Mainardi-Book ; Mainardi-C .
Historical Perspective.
While the modern definitions above were systematized in the 20th century, their conceptual roots lie in 19th-century problems. Abel’s 1823–1826 work on the tautochrone implicitly used what we now call the Caputo derivative, and his solution method anticipated the Riemann–Liouville fractional integral. Thus, the tautochrone problem not only motivated the first concrete application of fractional operators but also shaped their early development.
V Concluding Remarks